(10pts) 2. Assume X is normally distributed with a mean of 5 and variance of 16. Determine the value of x that solves each of the following: P(x < X < 9) = 0.2 b) P(-x < X-5<x) = 0.99
Answer #1

a. P(x < X < 9)

\Rightarrow P(\frac{x-5}{\sqrt{16}}<\frac{X-5}{\sqrt{16}}<\frac{9-5}{\sqrt{16}} )

\Rightarrow P(\frac{x-5}{4}<Z<1 ) = 0.2

\Rightarrow P(Z<1 )- P(Z<\frac{x-5}{4}) =0.841 -P(Z<\frac{x-5}{4})= 0.2

\Rightarrow P(Z<\frac{x-5}{4})= 0.841-0.2 = 0.641

we know,

P(Z<0.361) = 0.641

So, (x-5 / 4) = 0.641

So, x = 7.564

b. P(-x<X-5<x) = 0.99

\Rightarrow P(-x/\sqrt{16}<\frac{X-5}{\sqrt{16}}<x/\sqrt{16}) = 0.99

\Rightarrow P(-x/4<Z<x/4) = 0.99

We know, P(-2.576 < Z < 2.576) = 0.99

So, (x/4) = 2.576

or, x = 10.304